OpenStudy (goformit100):

A is 3 years older to B and 3 years younger to C, while B and D are twins. How many years older is C to D? A. 2 B. 3 C. 6 D. 12

OpenStudy (anonymous):

\[ A=B+3=C-3\\ B = D = C-6 \]

OpenStudy (goformit100):

thanks